### v2

# Learning Smooth State-Dependent Traversability from Dense Point Clouds

Abstract
A key open challenge in off-road autonomy is that the traversability of terrain often depends on the vehicle’s state. In particular, some obstacles are only traversable from some orientations. However, learning this interaction by encoding the angle of approach as a model input demands a large and diverse training dataset and is computationally inefficient during planning due to repeated model inference. To address these challenges, we present SPARTA, a method for estimating approach angle conditioned traversability from point clouds. Specifically, we impose geometric structure into our network by outputting a smooth analytical function over the 1-Sphere that predicts risk distribution for any angle of approach with minimal overhead and can be reused for subsequent queries. The function is composed of Fourier basis functions, which has important advantages for generalization due to their periodic nature and smoothness. We demonstrate SPARTA both in a high-fidelity simulation platform, where our model achieves a 91% success rate crossing a 40m boulder field (compared to 73% for the baseline), and on hardware, illustrating the generalization ability of the model to real-world settings. Our code will be available at https://github.com/neu-autonomy/SPARTA.
Keywords: Off-road Autonomy, Traversability, Geometric Deep Learning
1 Introduction
Figure 1: Dense point cloud captures geometric detail of the environment (Top right), while discretizing it into elevation map results in information loss. The underlying terrain is clearly distinguishable from noise (leaves, sensor noise) in the point cloud, but doing so is hard in an elevation map, leading to wrong risk estimation (Viridis Circle where Green denotes low risk).
Many applications for autonomous mobile robots involve traversing rough, off-road environments [1, 2, 3]. One key to reliably navigate in these environments is to estimate terrain traversability, i.e., the ability to understand the interaction between the robot and the terrain, and assess relevant risk metrics (e.g., potential vehicle damage). Despite recent success in identifying traversable terrain [4, 5, 6, 7], motion planning in geometrically complex environments (e.g., dense forests, boulder fields) remains a challenge. Specifically, the risk associated with traversing a given terrain is often a complicated function of the robot’s state, for example its speed and angle of approach, and the terrain geometry [^1] We focus on angle-dependency in this paper, and leave the extension to other state variables to future work. . It is especially unclear how to account for the angle of approach (e.g., using hand-crafted risk heuristics based on terrain elevation [8, 9, 10]), as this requires both detailed terrain information and nuanced consideration of the underlying geometry.
In contrast, learning-based approaches have promise in being able to flexibly model the complex relationship between terrain and traversal risk. In particular, point cloud-based approaches are suitable for the problem setting, because point clouds are capable of capturing fine geometric details of the environment that would be missed in other representations like an elevation map. As shown in Fig. 1, in the point cloud, the terrain (ground) is distinguishable from noise (e.g. leaves). As a result, a point cloud-based method would prefer the safe approach angle (Red) over the risky one (Magenta) due to a relatively smooth elevation change. However, noise and terrain cannot be easily separated in the elevation map despite its high resolution (5cm). The terrain appears smoother from the risky angle (Magenta) thus an elevation map-based method would (incorrectly) select it instead.
In this paper, we propose a technique to learn risk variables of interest, based on a point cloud representation of the terrain. Specifically, to model system noise implicitly captured in the dataset, we represent risk as a categorical distribution due to its ability to approximate arbitrary distributions [4]. However, a point cloud-based network that naively encodes the angle of approach would be hard to train in practice due to the sparsity of data in robotics applications. As a result, our model takes in the point cloud and predicts an analytical function (represented using Fourier basis functions) that maps angle of approach to the target risk variable distribution. The choice of Fourier basis ensures that the resulting function is periodic, thus avoiding the wrap-around discontinuity [11], and has important advantages for generalization in the training process due to its upper-bounded Lipschitz constant by construction. The function, once computed for a terrain patch, can be queried efficiently with minimal computation overhead for any angle of approach at any timestep when integrated with modern planners like MPPI [12, 13]. To summarize, in this paper we propose SPARTA (Smooth Point-cloud Approach-angle Reasoning for Terrain Assessment) with the following key contributions:
  • A novel learning-based approach leveraging Fourier basis functions, which estimates angle-of-approach dependent traversal risk for complex terrains represented as point clouds.
  • Theoretical analysis of the advantage of imposing smoothness by using Fourier basis functions, with connection to Lipschitz continuity and model generalization capability.
  • Evaluations of the proposed pipeline through learning to predict potential vehicle damage, a risk variable prior works handle using handcrafted rules, with extensive results both in a high-fidelity simulator [14] and on hardware, with integration into a MPPI planner [12]. In Appendix G we discuss SPARTA’s generalization to other problem settings.
2 Related Work
Traversability Estimation
Traversability estimation aims to identify terrain suitable for robot navigation. Traversability estimates typically rely on either manually designed features [8, 10] or are derived from learned vehicle movement patterns [15, 4, 6, 7, 5]. Both [8] and [10] use local elevation measurements to estimate stepping difficulty using heuristics such as step-wise elevation change. On the other hand, learned methods usually rely on multiple sensing modalities, especially elevation maps and semantic maps [15, 4, 6, 7, 5]. WVN [5] uses camera input and trains a traversable terrain classifier online leveraging pretrained features from DINO-ViT [16]. Meanwhile, [6, 7] fuse semantic and elevation measurements, and use separate decoder heads to predict dense semantic and elevation maps, which are used to score planner rollouts using various heuristics. [15, 4] use semantic and elevation maps to predict traversability distributions represented as linear/angular traction and incorporate the worst-case traction into the vehicle dynamics model for risk-aware planning. Several works modeled the state-dependency of traversability by encoding robot state as part of the neural networks’ input [17, 18, 19]. However, there performances suffer from data sparsity and the dimensionality mismatch between exteroceptive inputs and low dimensional state space.
Learning Harmonic Functions
Geometric deep learning has found its place in numerous machine learning applications. For example, FBM [20] uses Fourier basis functions to composite long-time series data and achieved state-of-the-art performance on multiple benchmarks. NeuRBF [21] replaces traditional grid-based structures with radial basis functions for constructing implicit neural representation of scenes, allowing the model to better adapt to the underlying signal structures. In robotics, OrbitGrasp [22] leverages spherical harmonics functions to represent a grasping quality function. Off-road autonomy, especially traversability estimation, is a setting with rich geometric structure, and as a result we propose to leverage Fourier basis functions to exploit this intrinsic prior.
Placement of This Work
Unlike previous traversability estimation algorithms, SPARTA explicitly models the angle dependency of risk variables by exploiting the geometric structure of angle of approach (the 1-Sphere). Because we impose few assumptions on the risk variable of interest (i.e., it has clear lower and upper bounds and does not change abruptly with a small change in angle of approach), our approach could be used to extend (most) existing traversability estimation methods for more accurate and granular prediction. Besides tire deformation (demonstrated in our experiments Section 4), we show examples of variables with a similar geometric structure in Appendix G.
3 Method
Figure 2: Overall pipeline of the proposed algorithm. For all terrain patches we are interested in, if its Fourier coefficients are not in the database, we compute them and store in the database. Otherwise we retrieve the Fourier coefficients from the database and construct the analytical functions, which are queried to compute the risk variable distribution, whose tail distribution is considered for risk aware planning.
In this section we introduce SPARTA, a method for predicting the risk associated with traversing a terrain patch at a given angle based on prior data traversing a wide variety of terrains (represented as dense point cloud). One challenge in traversability estimation is handling aleatoric uncertainty, the inherent uncertainty in a system which cannot be reduced with more data [23]. SPARTA captures this uncertainty by treating risk as a random variable – a categorical distribution with $B$ discrete bins – whose concentration parameters are predicted by our model (Section 3.1 and Section 3.2). The tail of the distribution is used to estimate risk for risk-aware planning (Section 3.4).
The primary novelty in SPARTA is in representing this risk distribution. SPARTA predicts an analytical function for each bin, and each function smoothly maps any angle of approach to the corresponding risk concentration parameter (Section 3.2). We characterize the smoothness of these risk functions (Section 3.3), which we hypothesize to be useful for generalization across approach angles (empirically supported in Section 4.2). Once computed, the function can be re-used for subsequent queries for the same terrain patch, thus avoiding repeated neural network inference during planning. We include an experiment demonstrating the query efficiency of our model in Appendix C. An overview of the risk estimation pipeline and its application to planning is shown in Fig. 2.
3.1 Modeling Aleatoric Uncertainty with Categorical Distributions
3.2 Leveraging $S^{1}$ as a Continuous Representation of Approach Angle
3.3 Imposed Smoothness Improves Prediction Generalization
3.4 Incorporating the Learned Risk Model into a Motion Planner
4 Experiments
In this section, we demonstrate the application of the proposed framework to assess potential vehicle damage (e.g., flat tires, broken axle). Although vehicle damage frequently occurs during real off-road autonomous operations, previous research has largely overlooked this risk metric or relied on relatively simple handcrafted heuristics (such as stepwise elevation change) to model it [8, 9, 10].
4.1 Tire Deformation As A Surrogate Metric
4.2 Model Performance
4.3 Planning in Simulation
4.4 Planning on Hardware
5 Conclusion
In this paper, we propose SPARTA, a general framework for estimating terrain traversal risk conditioned on vehicle angle of approach. Specifically, we identify and exploit the geometric structure in the problem setting, and propose to predict a smooth analytical function over the 1-Sphere to achieve better data efficiency and generalization performance. During planning, the function is queried to construct a categorical distribution for a risk variable of interest at any angle of approach, and the distribution’s tail risk (CVaR) is considered to account for aleatoric uncertainty. Experiments in simulation and on hardware demonstrate the effectiveness of the proposed approach. Most importantly, off-road problems are naturally rich in geometric structure, and our formulation is general enough such that it finds application in numerous different problem settings. Future work may further explore the geometric structure in other off-road autonomy problems.
6 Limitation
Our work has several limitations. For example, in this work we assume the environment is fully observable, i.e. a prebuilt point cloud map is accessible during planning. While the point cloud map we can build on the fly is sufficiently dense, the model’s performance without a prebuilt map is still to be tested. A potential solution to this problem is to penalize terrain with limited observation during planning or incorporate depth inpainting as in [7, 6]. In addition, this work focused on estimating traversability given geometric information, whereas semantic information is also important to consider in many off-road settings. We anticipate that future work could incorporate semantic information by performing sensor fusion with point pillar feature in birds-eye-view [7]. While the smoothness assumption in Section 3.3 is reasonable, settings that potentially break this assumption might exist in the real world. Quantitative analysis to identify such failure modes can be a valuable addition to our framework. Last but not least, tire deformation is a metric that’s only available in simulation. However, we are not interested in predicting tire deformation in the real world, but rather using it as a surrogate metric for terrain property (i.e. how risky is the terrain). Future work could explore other variables to quantify vehicle damage.
Acknowledgments
This research was sponsored by the DEVCOM Army Research Laboratory (ARL) under SARA CRA W911NF-24-2-0017. Distribution Statement A: Approved for public release; distribution is unlimited.
Appendix A Derivation of Lipschitz Constant
With a slight abuse of notation $F$, LABEL:eqn:fourier can be written in the following way:
$G(\phi)$ $=\frac{a_{i}^{0}}{2}+\sum_{k=1}^{n}[a_{i}^{k}\cos(k\phi)+b_{i}^{k}\sin(k\phi)]$ (1)
$F(\phi)$ $=\sigma(G(\phi))$ (2)
where $\sigma(\cdot)$ is the sigmoid function and $a_{i}^{0}$, $a_{i}^{k}$, and $b_{i}^{k}$ are the Fourier coefficients predicted by our network.
We will derive an upper bound on the Lipschitz constant of $F(\phi)$, via the chain rule:
$|F^{\prime}(\phi)|$ $=|\sigma^{\prime}(G(\phi))|\cdot|G^{\prime}(\phi)|.$ (3)
First we differentiate $G(\phi)$ w.r.t. $\phi$:
$G^{\prime}(\phi)$ $=\sum_{k=1}^{n}ka_{i}^{k}\cos(k\phi)-kb_{i}^{k}\sin(k\phi)$ (4)
$=\sum_{k=1}^{n}k[a_{i}^{k}\cos(k\phi)-b_{i}^{k}\sin(k\phi)].$ (5)
For every $k\in[1,n]$, let:
$R_{k}$ $=\sqrt{(a_{i}^{k})^{2}+(b_{i}^{k})^{2}}$ (6)
$\Delta_{k}$ $=\arctan(\frac{-b_{i}^{k}}{a_{i}^{k}}),$ (7)
Then Eq. 5 can be rewritten as a phase shift leveraging the cosine addition formula:
$G^{\prime}(\phi)$ $=\sum_{k=1}^{n}k[R_{k}\cos(k\phi)\frac{a_{i}^{k}}{R_{k}}-R_{k}\sin(k\phi)\frac{b_{i}^{k}}{R_{k}}]$ (8)
$=\sum_{k=1}^{n}k[R_{k}\cos(k\phi)\cos(\Delta_{k})+R_{k}\sin(k\phi)\sin(\Delta_{k})]$ (9)
$=\sum_{k=1}^{n}kR_{k}\cos(k\phi-\Delta_{k}).$ (10)
Notice that the maximum value of a $\cos(\cdot)$ is 1, as a result, Eq. 10 is upper bounded by:
$G^{\prime}(\phi)$ $\leq\sum_{k=1}^{n}kR_{k}$ (11)
$=\sum_{k=1}^{n}k\sqrt{(a_{i}^{k})^{2}+(b_{i}^{k})^{2}}.$ (12)
Because the sigmoid function is Lipschitz with a Lipschitz constant of $\frac{1}{4}$, by the chain rule:
$|F^{\prime}(\phi)|$ $=|\sigma^{\prime}(G(\phi))|\cdot|G^{\prime}(\phi)|$ (13)
$\leq\frac{1}{4}\sum_{k=1}^{n}k\sqrt{(a_{i}^{k})^{2}+(b_{i}^{k})^{2}}.$ (14)
As a result the Lipschitz Constant of $F(\phi)$ is bounded by:
$L^{i}\leq\frac{1}{4}\sum_{k=1}^{n}k\sqrt{(a_{i}^{k})^{2}+(b_{i}^{k})^{2}}.$
Note that the peak of the derivative of $G(\phi)$ and the sigmoid function $\sigma(\cdot)$ may not coincide, and as a result we expect the Lipschitz constant achieved in practice to be smaller.
Appendix B Model Architecture
All the learned models described in this paper (Ours, AngleInput, and AngleFree) predict categorical distributions composed of $B=8$ bins, parameterized by the concentration parameters $\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$.
All models share the same Point Pillars encoder architecture [30], which extracts a feature representation from the input point cloud $\mathbf{Q}$. Specifically, the point cloud $\mathbf{Q}$ is first uniformly scaled to fit within a cube defined by $x\in[-0.5,0.5]$, $y\in[-0.5,0.5]$, and $z\in[0,0.5]$. This normalized point cloud is then discretized into uniformly spaced pillars in the x-y plane, each with a grid size of $0.1m$. Points within each pillar are augmented with $x_{c},y_{c},z_{c},x_{p}$, and $y_{p}$ where the $c$ subscript denotes distance to the arithmetic mean of all points in the pillar and the $p$ subscript denotes the offset from the pillar’s $x,y$ center [30]. Notice we exclude the reflectance feature, yielding an 8-dimensional augmented input representation for each lidar point. For each pillar, the encoder samples up to 32 points and generates a corresponding 32-dimensional feature vector. The resulting output from the Point Pillars encoder is a pseudo-image tensor with dimensions $10\times 10\times 32$.
Next, we detail the remaining components of the model architecture.
B.1 Neural Network Used in SPARTA
The pseudo-image from the Point Pillars encoder [30] is subsequently processed by a two-layer convolutional backbone, specified as follows:
  • Input Channels: [32, 64]
  • Output Channels: [64, 256]
  • Kernel Size: [3, 3]
  • Stride: [2, 2]
  • Padding: [1, 1]
Each convolutional layer in the backbone employs batch normalization and ReLU activation. The resulting output tensor, of dimensions $3\times 3\times 256$, is passed through a max pooling operation to yield a compact 256-dimensional point cloud feature representation.
To represent the smooth analytical mapping $\phi\rightarrow\gamma_{\phi}^{i}$, we employ Fourier basis functions up to a maximum frequency of $n=3$, specifically the basis set $\{1,\cos(\phi),\sin(\phi),\cos(2\phi),\sin(2\phi),\cos(3\phi),\sin(3\phi)\}$. The choice of $n$ is justified in Appendix E with an ablation study. Consequently, for each bin $i\in[1,8]$, the network predicts 7 Fourier coefficients. Thus, the SPARTA decoder consists of a multilayer perceptron (MLP) that transforms the 256-dimensional point cloud features into a 56-dimensional output vector $\mathbf{F}^{1},\dots,\mathbf{F}^{8}]$. This MLP decoder contains one hidden layer with 512 neurons and utilizes layer normalization and ReLU activation functions.
B.2 AngleInput
The pseudo-image from the Point Pillars encoder [30] is subsequently processed by a three-layer convolutional backbone, specified as follows:
  • Input Channels: [32, 64, 128]
  • Output Channels: [64, 128, 256]
  • Kernel Size: [3, 3, 3]
  • Stride: [2, 2, 2]
  • Padding: [1, 1, 0]
The encoder for angle of approach $\phi$ is a MLP that maps 1-dimensional input $\phi$ to a 32-dimensional embedding with one hidden layer with 16 neurons. The decoder merges the point cloud feature (256-dimensional) and the angle feature (32-dimensional), and outputs the 8 concentration parameters $\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$, with a hidden layer with 64 neurons. Both MLPs utilizes layer normalization and ReLU activation functions.
B.3 AngleFree
The AngleFree baseline network uses the same convolutional backbone as AngleInput. A decoder head transforms the 256-dimensional point cloud feature to the 8 concentration parameters $\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$, with two hidden layers with [128, 64] neurons.
Appendix C Inference Efficiency
Figure 3: Runtime of AngleInput (Green) increases with model complexity, while the runtime of ours (Red) remains roughly constant. This is because during planning, we only need to perform a low dimensional dot product and a sigmoid pass during inference, whereas AngleInput requires repeated neural network inference. The x-axis denotes additional hidden layers in the decoder and additional frequencies in our model, i.e., $n\in[3,13]$.
To demonstrate the query efficiency of our model compared to AngleInput, we measure their respective inference runtimes. Specifically, for AngleInput, we measure the runtime of encoding the approach angle, decoding the concentration parameters, and computing the categorical distribution (LABEL:eqn:PMF). For our model, we measure the runtime associated only with computing the concentration parameters (LABEL:eqn:fourier) and the categorical distribution (LABEL:eqn:PMF). Note other components (e.g. extracting feature vector from point cloud) of both models can be precomputed or reused. Considering that more complex point clouds require higher-dimensional point cloud features and deeper network architectures, we perform a stress test by progressively increasing the number of layers in both models, as well as the maximum frequency $n$ in our model. Both models process 25,000 queries with average runtimes reported in Fig. 3. Our model demonstrates an order-of-magnitude faster runtime compared to AngleInput. Most importantly, the runtime of our model remains constant despite increased complexity, as it only involves a single batched, low-dimensional dot product and sigmoid computation  [^2] The dimension of the dot product depends only on the choice of maximum frequency $n$. In fact the dimension is exactly $2n+1$.. In contrast, the runtime of AngleInput increases with complexity.
Appendix D Data Collection Detail
In this section we provide more details of our data collection process using the automotive simulator BeamNG.tech [14]. We intend to generate random obstacles for the vehicle to drive over, and collect the tire deformation during the traversal. We choose a patch size of $1\times 1m$ that covers the footprint of the vehicle wheel, and randomly place (non-overlapping) obstacles in the patch. Maximum elevation of the obstacle is set to $0.5m$ (the vehicle’s minimum ground clearance). For every patch, we collect its point cloud and downsample to a maximum of 1024 points. To minimize the influence of vehicle dynamics (for the purpose of generalization to unseen hardware), we place the vehicle such that only one wheel contacts the obstacle. The deformation extent $d_{w}=\frac{r_{w}-r_{\text{inner}}}{r_{\text{outer}}-r_{\text{inner}}}$ of all contacting tire node $w$ are recorded during the traversal. We discard all samples below 0.2 (the tire deformation when the vehicle is static on flat ground) and divide the rest into a histogram with 8 bins, which forms the empirical distribution $\mathbf{y}$. In total, 24,000 samples $\{\mathbf{Q}_{k},\mathbf{y}_{k},\phi_{k}\}$ are collected. We augment the samples by applying 8 random rotations (to both the point cloud and the angle of approach), resulting in a total of 216,000 samples. We further add random perturbation to the point clouds for more robust performance. We use 90% of the data for training and 10% for testing.
Appendix E Traversing a Row of Obstacles with Ablation Study
Figure 4: (a): Overview of the test environment and the vehicle. (b)-(d): Example point cloud (colored by elevation) and the CVaR (Viridis arces around the point clouds) predicted by our model at the candidate angles of approach.
Algorithm $\alpha$ $n$ Suc. $\uparrow$ Dmg. $\downarrow$
Ours 0.9 3 95 5
0.9 1 91 9
0.9 5 94 6
0 3 89 11
AngleInput 0.9 92 8
0 89 11
Elev 73 27
Table 1: Number of success (Suc.) and vehicle damage (dmg.) in simulation. Our model outperforms AngleInput and Elev. The performance of the models decrease when we ablate CVaR with mean. Best , Second best.
In this experiment, we test our model’s ability to identify safe traversal points and angles of approach in a row of packed, randomly placed obstacles. An overview of the environment, a few example point clouds (colored by elevation), and the CVaRs predicted by our model are shown in Fig. 4. This environment is simpler than the boulder field in Section 4.3 and thus helpful in isolating the influence of the variables we are interested in. We consider two baselines, the AngleInput and a heuristic baseline that chooses the point with lowest maximum elevation under vehicle’s wheel footprint (Elev). Since the obstacles are placed on flat ground, the Elev baseline is equivalent to finding the plan with lowest stepping difficulty as in [8, 10]. As shown in Table 1, our model achieves higher success rate (S.R., the vehicle drives over the obstacle without damage) than the AngleInput, but both learning-based method significantly outperforms Elev. This is expected because the obstacles are spawned with similar sizes, and as a result Elev degenerates to random selection, leading to much lower success rate and thus high risk of vehicle damage.
Ablating CVaR
To demonstrate the effectiveness of using CVaR during planning to capture worst-case risk, we ablate it by setting $\alpha=0$, which is equivalent to using the mean of the predicted distribution. As reported in Table 1, the performance of both our model and AngleInput dropped. This is because accounting for worst-case (tail) risk is crucial in safety-critical tasks, and CVaR is a principled way for representing tail risk in robotics [28].
Ablating Maximum Frequency
Figure 5: Test $\text{EMD}^{2}$ loss of our model with varying maximum frequency. They achieve similar test loss.
As demonstrated in Section 3.2 and Section 3.3, the choice of maximum frequency $n$ is closely related to the smoothness and generalization of the mapping from angle of approach $\phi$ to each of the concentration parameters $\gamma_{\phi}^{i}$. Therefore, we adjust the number of Fourier bases $n$ in LABEL:eqn:fourier and analyze the results in this section. We consider the cases $n=1$ and $n=5$, and visualize the test $\text{EMD}^{2}$ loss in Fig. 5. Overall, we observe similar test $\text{EMD}^{2}$ loss for all three models. In Table 1 we include their performances on the aforementioned experiment, on which they achieve success rate 91% and 94%, respectively. Notice the model with $n=1$ achieves lower success rate compared to the other two models. We hypothesize this is because the basis functions can be too simple to capture the change of $\gamma_{\phi}^{i}$ with $\phi$, i.e. it might not be expressive enough to parameterize the mapping we are interested in. Although the model with maximum frequency $n=5$ achieves similar success rate as ours ($n=3$), an overly high maximum frequency $n$ introduces higher upper bound on the Lipschitz constant of the mapping and can introduce more abrupt changes in $\gamma_{\phi}^{i}$ with small changes in $\phi$, as explained in Section 3.3. Therefore, to balance expressivity and generalization, we choose $n=3$ in our model and our experiments. However, $n$ is an easily changed hyperparameter that can be adjusted depending on the frequency composition of the mapping function we are trying to approximate.
Appendix F Hardware Specification
Our robot is built on top of the AgileX Scout Mini chassis, a skid-steering wheeled platform. It has a footprint of approximately $0.6\times 0.6m$ and can achieve a maximum speed of 3 m/s. The robot is equipped with an Ouster OS1 LiDAR with 128 channels. The LiDAR runs with a horizontal resolution of 1024 at 10 Hz, and we use its internal 6-axis IMU (100 Hz). The point cloud and IMU measurements are processed using DLIO [29] for prebuilding the point cloud map and online localization. The onboard PC is an ASUS ROG NUC 970, with an Intel Core Ultra 9 CPU (16-Core), 32 GB RAM, and a Nvidia GeForce RTX 4070 GPU (8 GB).
Appendix G Generalization to Other Problem Settings
Figure 6: (a): Obstacle and directions of the traversals. (b): The linear traction distribution of the forward (Blue) and backward (Orange) traversal. The backward distribution has significantly more probability mass in the lower bins. An algorithm not modeling the angle-dependency of linear traction (Green) would not be able to identify the obstacle as traversable from the forward orientation.
In this section, we first demonstrate other problem settings that could benefit from modeling angle-dependency. We use the angle-dependency of linear traction (the robot’s ability to execute commanded velocity on the given terrain [4]) as an example, and show how our approach can be applied to this problem setting. To construct an empirical dataset, we command the vehicle to drive at a constant linear velocity of 0.5 m/s over the obstacle, along the forward (Blue) and the backward (Orange) direction shown in Fig. 6(a). We record the achieved linear velocity (estimated via DLIO [29]), and compute the linear traction as the ratio between the achieved linear velocity and the commanded linear velocity (0.5 m/s). The data points are discretized into histograms and then normalized into categorical distributions. As shown in Fig. 6(b), the backward traversal (Orange) achieved significantly lower linear tractions compared to the forward traversal (Blue). This matches our intuition because the obstacle is smooth from the forward direction and sharp from the other. A system that models the angle-dependency of linear traction would thus be able to capture this difference. Notice SPARTA can be applied to this problem setting with minimal changes (e.g. adjusting input point cloud size) using an empirical dataset of linear traction distributions. However, systems that do not model angle-dependency (Green) [4] could be overly conservative and never command the robot to traverse the obstacle from the forward direction, which is traversable in practice.
In fact, we anticipate any variables that exhibit smooth and continuous behaviors, especially those that depends on variables with Lie group structure, could similarly benefit from our representation framework. SPARTA can be easily modified to use their respective basis functions and predict the corresponding coefficients.
Appendix H Example CVaR Predictions of SPARTA
Figure 7: Example predictions of SPARTA. Rainbow denotes point elevation, and Viridis circles around the point clouds denote the CVaR at the corresponding angle of approach. The CVaR matches our intuition because it indicates the obstacles are more risky from orientations with sharp edges and less risky from orientations with smooth elevation change.
In Fig. 7, we visualize some example obstacles in our test set and the CVaR predicted by our model with angle of approach $\phi\in\{0,1,\dots,359\}$. The Rainbow color map denotes the elevation of the point (Red denotes high elevation), and Viridis denotes the predicted CVaR (Yellow corresponds to high CVaR). A trend we notice from the examples is that our model predicts lower CVaR for angles where the obstacle has a smoother elevation change. This is aligned with our intuition, further supporting the validity of our approach for identifying risky obstacles and angles of approach.